OpenStudy (anonymous):

if f(x)=x^3 & g(x)=x-1, which of the following statements is true?

OpenStudy (anonymous):

A) (gof)(0)=3 B) (fog)(0)=1 C) (fog)(2)=1 D) (gof)(2)=6

OpenStudy (anonymous):

Can you please check the question again? Just once.

OpenStudy (anonymous):

yeah, its all correct

OpenStudy (anonymous):

Okay hold on.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

if f(x)=x^3 & g(x)=x-1, which of the following statements is true? F(x) = x^3 G(x) = x -1 1. F[G(x)] = FoG(0) = F[G(0)] = F[0 - 1] = F[-1] = -1 2. GoF(0) = G[F(0)] = G[0] = -1 3. FoG(2) = F[G(2)] = F[1] = 1 Third choice C.

OpenStudy (anonymous):

f(gx)=(x-1)^3,, f(2)=(2-1)^3 ,, C